# ComKD-CLIP: Comprehensive Knowledge Distillation for Contrastive Language-Image Pre-traning Model

### v1

@@ Abstract
Contrastive Language-Image Pre-training (CLIP) excels in integrating semantic information between images and text through contrastive learning techniques. It has achieved remarkable performance in various multimodal tasks. However, the deployment of large CLIP models is hindered in resource-limited environments, while smaller models frequently fall short of meeting performance benchmarks necessary for practical applications. In this paper, we propose a novel approach, coined as ComKD-CLIP: Comprehensive Knowledge Distillation for Contrastive Language-Image Pre-traning Model, which aims to comprehensively distill the knowledge from a large teacher CLIP model into a smaller student model, ensuring comparable performance with significantly reduced parameters. ComKD-CLIP is composed of two key mechanisms: Image Feature Alignment (IFAlign) and Educational Attention (EduAttention). IFAlign makes the image features extracted by the student model closely match those extracted by the teacher model, enabling the student to learn teacher’s knowledge of extracting image features. EduAttention explores the cross-relationships between text features extracted by the teacher model and image features extracted by the student model, enabling the student model to learn how the teacher model integrates text-image features. In addition, ComKD-CLIP can refine the knowledge distilled from IFAlign and EduAttention leveraging the results of text-image feature fusion by the teacher model, ensuring student model accurately absorbs the knowledge of teacher model. Extensive experiments conducted on 11 datasets have demonstrated the superiority of the proposed method.
@@ Introduction
Figure 1: (a) The schematic of proposed ComKD-CLIP framework. (b) Comparison of the Harmonic Mean (HM) for generalization from base to novel categories. All methods utilize the ViT-B/16 image encoder from the pretrained CLIP model. ComKD-CLIP is proud to achieve best performance in 8 out of 11 diverse recognition datasets.
@@
Contrastive Language-Image Pre-training (CLIP) (Radford et al. 2021) has emerged as a leading pre-trained Visual Language Models (VLMs) due to its ability to effectively learn the fusion of semantic information between images and text through contrastive learning. This ability has enabled CLIP to demonstrate strong performance across various multimodal tasks such as image recognition, visual question answering, and image description generation (Li et al. 2023a; Wei et al. 2022; Singh et al. 2022). However, large CLIP models cannot be deployed in resource-constrained environments and the performance of small CLIP models often falls short of application requirements. These limitations significantly restrict the practical applicability of CLIP models.
@@
Knowledge Distillation (KD) (Hinton, Vinyals, and Dean 2015) has been introduced CLIP to address these issues. CLIP-TD (Wang et al. 2022) distills knowledge from both CLIP’s vision and language branches into existing architectures for VL tasks. TinyCLIP (Wu et al. 2023) proposes affinity mimicking and weight inheritance to improve the performance of small models by leveraging large-scale models. CLIPPING (Pei et al. 2023) proposes a new layer-wise alignment with the student as the base, which enables the students to fully absorb the knowledge of the teacher. PromptKD (Li et al. 2024b) uses unlabeled domain data to perform prompt-based knowledge distillation for CLIP, which greatly improve the performance of small model. CLIP-KD (Yang et al. 2024) proposes several distillation strategies to examine the effectiveness of CLIP-KD. However, these methods distill the knowledge of the teacher model based on the results of text-image feature fusion, ignoring the knowledge embedded within the fusion process. This oversight substantially hinders the student model to absorb the teacher model’s knowledge.
@@
To ensure that the student model can fully absorb the knowledge of the teacher model, we endeavor to distill the knowledge embedded in the feature fusion process of a large model and refines the distilled knowledge leveraging the feature fusion results of the large model. Accordingly, we propose ComKD-CLIP: Comprehensive Knowledge Distillation for Contrastive Language-Image Pre-traning Model. ComKD-CLIP is composed of tow key modules: Image Feature Alignment (IFAlign) and Educational Attention (EduAttention). During the feature fusion stage: IFAlign ensures that the image features extracted by the student model closely match those extracted by the teacher model. This alignment enable the student to absorb the teacher’s knowledge of how to extract image feature. Concurrently, EduAttention explores the cross-relationship between the text features extracted by the teacher model and the image features extracted by the student model. Through this strategy, EduAttention enables the student model to comprehend and emulate the teacher model’s abilities for integrating image and text features, thus enriching its own mutlimodal understanding. Furthermore, to make the student model fully absorb the knowledge of the teacher model, ComKD-CLIP refines the knowledge distilled from ImgAlign and EduAttention leveraging the results of feature fusion. The schematic of proposed ComKD-CLIP is shown as Fig. 1 (a). We also compare the performance of the proposed method with some state-of-the-art methods on 11 datasets. The experimental results are shown in Fig. 1 (b). It is worth mentioning that the proposed method has best performance in 8 out of 11 diverse recognition datasets.
@@
  • We refines the knowledge distilled from IFAlign and EduAttention leveraging the results of feature fusion, which can prompt the small model to fully absorb the knowledge of large model. Extensive experiments conducted on 11 datasets have demonstrated the superiority of the proposed method.
@@
KD aims to enable the small student model to absorb knowledge from the large teacher model, thus achieving comparable performance to the large model. It has achieved remarkable success in numerous vision tasks, including image segmentation (Liu et al. 2019; Yang et al. 2022), object detection (Jia et al. 2024; Wang et al. 2024), and pose estimation (Li et al. 2021). Recently, many researchers have endeavored to introduce KD into CLIP, motivated by the pressing need to surmount the operational challenges faced by large CLIP models in resource-limited environments, alongside the performance shortfalls exhibited by smaller models in practical applications.(Laroudie et al. 2023). CLIPPING (Pei et al. 2023) introduces a novel layer-wise alignment strategy that takes the student model as the foundation, enabling the student model to effectively absorb the knowledge from the teacher model. PromptKD (Li et al. 2024b), in a distinct approach, capitalizes on unlabeled domain data to facilitate prompt-based knowledge distillation within the CLIP paradigm, significantly bolstering the performance of smaller CLIP models. TinyCLIP (Wu et al. 2023), similarly targeting CLIP distillation, achieves commendable results through the innovative application of affinity mimicking and weight inheritance techniques. However, prior studies predominantly concentrate on distilling the knowledge of teacher model based on feature fusion results, overlooking the intricate knowledge encapsulated within the feature fusion process. In stark contrast to existing distillation methodologies, our proposed approach uniquely distills the knowledge inherent to the text-image feature fusion process within large CLIP models. In addition, by refining the distilled knowledge leveraging the feature fusion results, we enhance the capability of student model to absorb the nuanced knowledge from teacher model, thereby precipitating a marked improvement in their performance.
@@
CLIP is one of the most commonly used VLMs, comprising independent image and text encoder branches. It aligns and fuses images with text to learn a joint multimodal embedding space. In the image encoding branch, a labeled visual recognition dataset $D=\{x_{j},y_{j}\}_{j=1}^{M}$ is used as input. Each image $x$ from dataset $D$ is processed by the image encoder $f_{I}$ to obtain normalized image features $u=f_{I}(x)/||f_{I}(x)||_{2}\in\mathbb{R}^{d}$. Corresponding to the image recognition dataset $D$ are $N$ class names $\text{c}=\{c_{i}\}_{i=1}^{N}$. In the text encoder branch, the input data are text descriptions $t_{i}$ generated from the template “a photo of a {$c_{i}$}”. Each $t_{i}$, after being encoded by the text encoder, yields normalized text features $w_{i}=f_{T}(t_{i})/||f_{T}(t_{i})||_{2}\in\mathbb{R}^{d}$, where $d$ is the dimension of the text features. The collection of all text features $W=[w_{1},w_{2},\ldots,w_{N}]\in\mathbb{R}^{N\times d}$ serves as the classification weight matrix. Based on this data, the classification output probability can be computed using Equation 1.
@@
KD originally proposed by Hinton et al. (Hinton, Vinyals, and Dean 2015), it transfers knowledge from a large, pretrained teacher model to a smaller, lightweight student model. The goal is for the student to absorb the teacher’s knowledge for efficient deployment. This process employs the KL divergence loss to align the feature distributions of both models. The KL divergence loss is defined as follows:
@@
Figure 2: Overview of our proposed ComKD-CLIP framework: (a) Utilization of the cue learning method with a well-trained large teacher CLIP model; (b) A smaller student CLIP model, which is to be trained with learnable cues and reuses the text features from the teacher model, requiring training only for the image encoder branch; (c) The schematic of IFAlign module; (d) The inference process within the student model, where the text encoder branch reuses the text features of the teacher model; (e) The schematic of EduAttention module.
@@
As illustrated in Fig. 2, our proposed ComKD-CLIP framework comprises two principal stages: the pretraining of the large CLIP teacher model and the subsequent training of a small CLIP student model. In the initial phase, as delineated in Fig.  2(a), the large CLIP teacher model is pretrained on a labeled domain dataset, $D_{labeled}=\{x_{i},y_{i}\}_{i=1}^{M}$, to enhance its performance, aligning with contemporary methodologies such as PromptSRC (Khattak et al. 2023b) and PromptKD (Li et al. 2024b). Innovatively, we incorporate learnable prompts into both the image and text encoder branches of the teacher model via a concatenation strategy. The image and text data from the labeled domain dataset are processed through the image encoder $f_{V}^{t}$ and text encoder $f_{T}^{t}$, respectively, producing image features $u^{p}_{t}\in\mathbb{R}^{d}$ and text features $w^{p}_{t}\in\mathbb{R}^{d}$. The ultimate output logits $q^{t}$ is calculated by matrix multiplication between $u^{p}_{t}$ and $w^{p}_{t}$. The training of the teacher model entails minimizing the cross-entropy loss between the predicted probability distributions and the true labels, thereby optimizing the model’s parameters. This rigorous pretraining phase ensures that the teacher model acquires robust knowledge that can be effectively distilled to the student model during the latter stage of our framework.
As depicted in Fig. 2(b), the student CLIP model directly capitalizes on the pre-trained text features from the teacher model, thereby significantly curtailing the training costs with the text encoder branch. Simultaneously, a lightweight CLIP image encoder branch is engineered within the student model to decrease resource costs while maintaining competitive performance. During the processing of input data from the unlabeled domain dataset $D_{unlabeled}$ by the student model’s image encode, we incorporate the IFAlign module. This module serves to align the student model’s image features $u^{p}_{s}\in\mathbb{R}^{d}$ with the teacher model’s image features $u^{p}_{t}\in\mathbb{R}^{d}$, thereby facilitating the student model to absorb the knowledge of how the teacher model extracts salient image features. Subsequent to the feature alignment, the EduAttention module is introduced to explore the cross-relationship between the image features extracted by student model and text features provided by teacher model. This exploration enables the student model to learn the nuanced strategies employed by the teacher model for integrating text-image feature. In addition, we employ the KL divergence to minimize the discrepancy between the logits produced by the teacher and student models. This optimization ensures that the knowledge distilled by the student model is refined and closely mirrors that of the teacher, thus enhancing the student’s understanding of the teacher’s knowledge.
@@
IFAlign
The schematic of IFAlign is illustrated in Fig. 2(c). To make the image features extracted by the student model closely match those extracted by the teacher model, we align the mean and variance statistics for extracted features. The process of calculate can be formulated as follows:
@@
EduAttention
This schematic of EduAttention is illustrated in Figure 1(e). In this module, an attention mechanism is leveraged to explore the cross-relationship between the image features extracted by student model and text features provided by teacher model, which can facilitate the the student model to learn the nuanced strategies employed by the teacher model for integrating text-image feature. The specific calculation process can be formulated as follows:
@@
CLIP 72.43 68.14 70.22
CoOp 76.47 67.88 71.92
CoCoOp 75.98 70.43 73.10
MaPLe 76.66 70.54 73.47
PromptSRC 77.60 70.73 74.01
PromptKD 80.83 74.66 77.62
ComKD-CLIP
$\Delta$ +– +– +–
@@
CLIP 69.34 74.22 71.70
CoOp 82.69 63.22 71.66
CoCoOp 80.47 71.69 75.83
MaPLe 82.28 75.14 78.55
PromptSRC 84.26 76.10 79.97
PromptKD 86.96 80.73 83.73
ComKD-CLIP —- —- —-
$\Delta$ +—- —– —–
Distilled Knowledge Refinement. After the student model absorbs the knowledge of how the teacher model extracts image features and combines text-image features, we try to refine the absorbed knowledge based on the feature fusion results of teacher model. As depicted in Fig. 2, we utilize the KL divergence to minimize the discrepancy between the feature distribution produced by the teacher and student model. The specific process can be formulated as follows:
@@
where $q^{t}$ and $q^{s}$ represent the logits predicted by the teacher and student models, respectively, which is calculated by the corresponding image features and text features using Equation 1. $\tau$ is the temperature parameter, which is used to adjust the smoothness of the probability distribution.
Finally, we combine the student model’s alignment loss $\mathcal{L}_{\text{stu}}$ with the feature distribution loss $\mathcal{L}_{\text{align}}$ as the loss function of small CLIP model, the specific loss formula is shown as follows:
$\mathcal{L}_{\text{final}}=\mathcal{L}_{\text{stu}}+\mathcal{L}_{\text{align}}.\quad$ (9)
@@
Following (Zhou et al. 2022a) (Khattak et al. 2023a) (Khattak et al. 2023b) (Li et al. 2024b), we partition training and testing datasets into base and novel classes. Our teacher model, pre-trained via the PromptSRC approach (Khattak et al. 2023b), guides the training of student models using an unlabeled set. Post-distillation, we assess the students’ performance on both class types against the test set, serving as a measure of methodological generalization within the dataset. As shown in Table 1, we compare the performance of our proposed ComKD-CLIP with recent state-of-the-art methods including CLIP (Radford et al. 2021), CoOp (Zhou et al. 2022b), CoCoOp (Zhou et al. 2022a), MaPLe (Khattak et al. 2023a), PromptSRC (Khattak et al. 2023b), PromptKD (Li et al. 2024b) on 11 recognition datasets. In comparison with this state-of-the-art works, ComKD-CLIP shows highly competitive results in all 11 datasets.
@@
Similar to PromptKD (Li et al. 2024b), our teacher model undergoes pre-training on ImageNet (Deng et al. 2009). Subsequently, we utilize the training set of unlabeled target datasets to train the student models. Their performance is then evaluated on the test set post-training, with no data-specific fine-tuning applied. We compare our cross-dataset performance with previous methods in Table 2, our proposed ComKD-CLIP outperforms some state-of-the-art methods on 8 out of 10 datasets, achieving an average improvement of 0.74% over previous methods.
@@
Table 2: We evaluate the performance of ComKD-CLIP with some state-of-the-art mothods in a cross-dataset benchmark. Utilizing our pipeline, we conduct unsupervised Aligned distillation with unlabeled domain data in a transductive setting. The source model is pretrained on ImageNet. ”ZSL” indicates that the evaluation is conducted in a Zero-Shot Learning setting. ComKD-CLIP outperforms on 8 out of 10 datasets.
@@
Table 3: Comparison of Com-CLIP within domain generalization contexts. The method is trained on ImageNet and subsequently evaluated on datasets with domain shifts.
@@
To begin with, we try to explore the effectiveness of IFAlign module and EduAttention module. Specifically, we remove the IFAlign module and EduAttention module from ComKD-CLIP respectively and test the corresponding experimental results. As shown in Table 4, removing IFAlign module will greatly reduce the performance of ComKD-CLIP. Similarly, removing the EduAttention module also decrease the model’s performance. It is worth noting that when both the IFAlign module and the EduAttention module are removed, ComKD-CLIP has the worst performance. These experimental results fully prove that IFAlign module and EduAttention module can effectively prompt the student model to absorb the knowledge of how the teacher model extracts image features and combines text-image feature.
@@
Subsequently, we try to use different alignment methods in IFAlign to find the best alignment strategy. We respectively align image features by mean ($\mathcal{L}_{\text{align\_mean}}$), variance ($\mathcal{L}_{\text{align\_var}}$), and mean + variance ($\mathcal{L}_{\text{align}}$). The experimental results are shown in Table 5. It can be clearly seen that ComKD-CLIP has the best performance when we use the mean + variance ($\mathcal{L}_{\text{align}}$) method to align image features in IFAlign. This mainly because that mean+variance can jointly promote the alignment from the perspectives of center position and discreteness, thereby helping the student model to learn how the teacher model extract image features.
@@
Finally, we try to explore the best strategy to refine the knowledge distilled by IFAlign module and EduAttention module. Specifically, we use KL divergence, L1 and MSE methods to refine the distilled knowledge. The corresponding performance of ComKD-CLIP are shown in Table 6. It can be seen that KL divergence can better refine the distilled knowledge and make ComKD-CLIP perform best. This may be because KL divergence can make the logits of the student model better approximate the logits of the teacher model, thereby better refining the knowledge distilled by the student model.

### v3

@@ Abstract
Contrastive Language-Image Pre-training (CLIP) models excel in integrating semantic information between images and texts through contrastive learning techniques. It has achieved remarkable performance in various multimodal tasks. However, the deployment of large CLIP models is hindered in resource-limited environments, while smaller models frequently fail to meet the performance benchmarks required for practical applications. In this paper, we propose a novel approach, ComKD-CLIP: Comprehensive Knowledge Distillation for Contrastive Language-Image Pre-traning Model, which aims to comprehensively distill the knowledge from a large teacher CLIP model into a smaller student model, ensuring comparable performance with significantly reduced parameters. ComKD-CLIP is composed of two key mechanisms: Image Feature Alignment (IFAlign) and Educational Attention (EduAttention). IFAlign makes the image features extracted by the student model closely match those extracted by the teacher model, enabling the student to learn teacher’s knowledge of extracting image features. EduAttention explores the cross-relationship between text features extracted by the teacher model and image features extracted by the student model, enabling the student model to learn how the teacher model integrates text-image features. In addition, ComKD-CLIP can refine the knowledge distilled from IFAlign and EduAttention by leveraging the text-image feature fusion results of the teacher model, ensuring the student model accurately absorbs the teacher’s knowledge. Extensive experiments conducted on 11 datasets have demonstrated the superiority of the proposed method.
@@ Introduction
Figure 1: (a) Schematic of proposed ComKD-CLIP framework. (b) Comparison of the Harmonic Mean (HM) for generalization from base to novel categories. All methods utilize the ViT-B/16 image encoder from the pretrained CLIP model. ComKD-CLIP is proud to achieve the best performance in 8 out of 11 diverse recognition datasets.
@@
Contrastive Language-Image Pre-training (CLIP) (Radford et al. 2021) has emerged as a prominent pre-trained Visual Language Model (VLM) due to its ability to effectively integrate the semantic information between images and texts through contrastive learning techniques. This ability has enabled CLIP to demonstrate strong performance across various multimodal tasks such as image recognition, visual question answering, and image description generation (Li et al. 2023a; Wei et al. 2022; Singh et al. 2022). However, large CLIP models cannot be deployed in resource-constrained environments and the performance of smaller CLIP models often falls short of application requirements. These limitations significantly restrict the practical applicability of CLIP models.
@@
Knowledge Distillation (KD) (Hinton, Vinyals, and Dean 2015) has been introduced into CLIP to address these issues. CLIP-TD (Wang et al. 2022) distills knowledge from both CLIP’s vision and language branches into existing architectures for VL tasks. TinyCLIP (Wu et al. 2023) proposes affinity mimicking and weight inheritance to improve the performance of small models by leveraging large-scale models. CLIPPING (Pei et al. 2023) proposes a new layer-wise alignment with the student as the base, which enables the student to fully absorb the knowledge of the teacher. PromptKD (Li et al. 2024b) uses unlabeled domain data to perform prompt-based knowledge distillation for CLIP, which can greatly improve the performance of small model. CLIP-KD (Yang et al. 2024) proposes several distillation strategies to examine the effectiveness of CLIP-KD. However, these methods distill the knowledge of the teacher model based on the results of text-image feature fusion, ignoring the knowledge embedded within the fusion process. This oversight substantially hinders the student model to absorb the teacher model’s knowledge.
@@
To ensure that the student model can fully absorb the knowledge of the teacher model, we endeavor to distill the knowledge embedded in the feature fusion process of teacher model and refine the distilled knowledge leveraging the feature fusion results of teacher model. Accordingly, we propose ComKD-CLIP: Comprehensive Knowledge Distillation for Contrastive Language-Image Pre-traning Model. ComKD-CLIP is composed of two key modules: Image Feature Alignment (IFAlign) and Educational Attention (EduAttention). During the feature fusion stage: IFAlign ensures that the image features extracted by the student model closely match those extracted by the teacher model. This alignment enable the student to absorb the teacher’s knowledge of how to extract image feature. Concurrently, EduAttention explores the cross-relationship between the text features extracted by the teacher model and the image features extracted by the student model. Through this strategy, EduAttention enables the student model to comprehend and emulate the teacher model’s abilities for integrating image and text features, thus enriching its own mutlimodal understanding. Furthermore, to make the student model accurately absorb the knowledge of the teacher model, ComKD-CLIP refines the knowledge distilled from ImgAlign and EduAttention leveraging the feature fusion results of the teacher model. The schematic of proposed ComKD-CLIP is shown as Fig. 1 (a). We also compare the performance of the proposed method with some state-of-the-art methods on 11 datasets. The experimental results are shown in Fig. 1 (b). It is worth mentioning that the proposed method has best performance in 8 out of 11 diverse recognition datasets.
@@
  • We refine the knowledge distilled from IFAlign and EduAttention leveraging the feature fusion results of the teacher model, which can prompt the student model to accurately absorb the teacher model’s knowledge. Extensive experiments conducted on 11 datasets have demonstrated the superiority of the proposed ComKD-CLIP.
@@
KD aims to enable the small student model to absorb knowledge from the large teacher model, thus achieving comparable performance to the large model. It has achieved remarkable success in numerous vision tasks, including image segmentation (Liu et al. 2019; Yang et al. 2022), object detection (Jia et al. 2024; Wang et al. 2024), and pose estimation (Li et al. 2021). Recently, many researchers have endeavored to introduce KD into CLIP, motivated by the pressing need to surmount the operational challenges faced by large CLIP models in resource-limited environments, alongside the performance shortfalls exhibited by smaller models in practical applications.(Laroudie et al. 2023). CLIPPING (Pei et al. 2023) introduces a novel layer-wise alignment strategy that takes the student model as the foundation, enabling the student model to effectively absorb the knowledge from the teacher model. PromptKD (Li et al. 2024b), in a distinct approach, capitalizes on unlabeled domain data to facilitate prompt-based knowledge distillation within the CLIP paradigm, significantly bolstering the performance of smaller CLIP models. TinyCLIP (Wu et al. 2023), similarly targeting CLIP distillation, achieves commendable results through the innovative application of affinity mimicking and weight inheritance techniques. However, prior studies predominantly concentrate on distilling the knowledge of teacher model based on feature fusion results, overlooking the intricate knowledge encapsulated within the feature fusion process. In stark contrast to existing distillation methodologies, the proposed method uniquely distills the knowledge inherent to the text-image feature fusion process within teacher models. In addition, by refining the distilled knowledge leveraging the feature fusion results, ComKD-CLIP enables the student model to accurately absorb the nuanced knowledge from teacher model, thereby precipitating a marked improvement in their performance.
@@
CLIP is one of the most commonly used VLMs, comprising independent image and text encoder branches. It aligns and fuses images with texts to learn a joint multimodal embedding space. In the image encoding branch, a labeled visual recognition dataset $D=\{x_{j},y_{j}\}_{j=1}^{M}$ is used as input. Each image $x$ from dataset $D$ is processed by the image encoder $f_{I}$ to obtain normalized image features $u=f_{I}(x)/||f_{I}(x)||_{2}\in\mathbb{R}^{d}$. Corresponding to the image recognition dataset $D$ has $N$ class names $\text{c}=\{c_{i}\}_{i=1}^{N}$. In the text encoder branch, the input data is text descriptions $t_{i}$ generated from the template “a photo of a {$c_{i}$}”. Each $t_{i}$, after being encoded by the text encoder, yields normalized text features $w_{i}=f_{T}(t_{i})/||f_{T}(t_{i})||_{2}\in\mathbb{R}^{d}$, where $d$ is the dimension of the text features. The collection of all text features $W=[w_{1},w_{2},\ldots,w_{N}]\in\mathbb{R}^{N\times d}$ serves as the classification weight matrix. Based on this data, the classification output probability can be calculated as follows:
@@
KD is originally proposed by Hinton et al. (Hinton, Vinyals, and Dean 2015), it transfers knowledge from a large, pre-trained teacher model to a smaller, lightweight student model. It can help the student to absorb the teacher’s knowledge for efficient deployment. This process employs the KL divergence loss to align the feature distributions of both models. The KL divergence loss is defined as follows:
@@
Figure 2: Overview of our proposed ComKD-CLIP framework. (a) Utilization of the cue learning method with a well-trained large teacher CLIP model; (b) A smaller student CLIP model, which is trained with learnable cues and reuses the text features from the teacher model, requiring training only for the image encoder branch; (c) The schematic of IFAlign module; (d) The inference process within the trained student model, where the text encoder branch reuses the text features of the teacher model; (e) The schematic of EduAttention module.
@@
As illustrated in Fig. 2, our proposed ComKD-CLIP framework comprises two principal stages: the pretraining of the large CLIP teacher model and the subsequent training of a small CLIP student model. In the initial phase, as delineated in Fig.  2(a), the large CLIP teacher model is pretrained on a labeled domain dataset, $D_{labeled}=\{x_{i},y_{i}\}_{i=1}^{M}$, to enhance its performance, aligning with contemporary methodologies such as PromptSRC (Khattak et al. 2023b) and PromptKD (Li et al. 2024b). Innovatively, we incorporate learnable prompts into both the image and text encoder branches of the teacher model via a concatenation strategy. The image and text data from the labeled domain dataset are processed through the image encoder $f_{V}^{t}$ and text encoder $f_{T}^{t}$, respectively, producing image features $u^{p}_{t}\in\mathbb{R}^{d}$ and text features $w^{p}_{t}\in\mathbb{R}^{d}$. The ultimate output logits $q^{t}$ is calculated by Eq. 1. The training of the teacher model entails minimizing the cross-entropy loss between the predicted probability distributions and the true labels, thereby optimizing the model’s parameters. This rigorous pretraining phase ensures the teacher model acquires robust knowledge that can be effectively distilled to the student model during the latter stage of our framework.
As depicted in Fig. 2(b), the student CLIP model directly capitalizes on the pre-trained text features from the teacher model, thereby significantly curtailing the training costs with the text encoder branch. Simultaneously, a lightweight CLIP image encoder branch is engineered within the student model to decrease resource costs while maintaining competitive performance. During the processing of input data from the unlabeled domain dataset $D_{unlabeled}$ by the student model’s image encode, we incorporate the IFAlign module. This module serves to align the student model’s image features $u^{p}_{s}\in\mathbb{R}^{d}$ with the teacher model’s image features $u^{p}_{t}\in\mathbb{R}^{d}$, thereby facilitating the student model to absorb the knowledge of how the teacher model extracts salient image features. Subsequent to the feature alignment, the EduAttention module is introduced to explore the cross-relationship between the image features extracted by student model and text features provided by teacher model. This exploration enables the student model to learn the nuanced strategies employed by the teacher model for integrating text-image feature. In addition, we employ the KL divergence to minimize the discrepancy between the logits produced by the teacher and student models. This optimization ensures that the knowledge distilled by the student model is refined and closely mirrors that of the teacher, thus enhancing the student model to accurately absorb the knowledge of teacher model. Finally, the inference process of trained student model is shown in Fig. 2(d).
@@
IFAlign: The schematic of IFAlign is illustrated in Fig. 2(c). To make the image features extracted by the student model closely match those extracted by the teacher model, we align the mean and variance statistics for extracted features. The calculation process can be formulated as follows:
@@
EduAttention: The schematic of EduAttention is illustrated in Fig. 2(e). In this module, an attention mechanism is leveraged to explore the cross-relationship between the image features extracted by student model and text features provided by teacher model, which can facilitate the student model to learn the nuanced strategies employed by the teacher model for integrating text-image feature. The specific calculation process can be formulated as follows:
@@
CLIP 69.34 74.22 71.70
CoOp 82.69 63.22 71.66
CoCoOp 80.47 71.69 75.83
MaPLe 82.28 75.14 78.55
PromptSRC 84.26 76.10 79.97
PromptKD 86.96 80.73 83.73
ComKD-CLIP 87.37 80.59 83.84
$\Delta$ +0.41 -0.14 +0.11
@@
CLIP 72.43 68.14 70.22
CoOp 76.47 67.88 71.92
CoCoOp 75.98 70.43 73.10
MaPLe 76.66 70.54 73.47
PromptSRC 77.60 70.73 74.01
PromptKD 80.83 74.66 77.62
ComKD-CLIP 76.53 71.02 73.67
$\Delta$4.30 -3.64 -3.95
Distilled Knowledge Refinement: After the student model absorbs the knowledge of how the teacher model extracts image features and combines text-image features, we try to refine the absorbed knowledge based on the feature fusion results of the teacher model. As depicted in Fig. 2, we utilize the KL divergence to minimize the discrepancy between the feature distribution produced by the teacher and student models. The specific process can be formulated as follows:
@@
where $q^{t}$ and $q^{s}$ represent the logits predicted by the teacher and student models, respectively, which is calculated by the corresponding image features and text features using Eq. 1. $\tau$ is the temperature parameter, which is used to adjust the smoothness of the probability distribution.
Finally, we combine the student model’s alignment loss $\mathcal{L}_{\text{stu}}$ with the feature distribution loss $\mathcal{L}_{\text{align}}$ as the final loss function to train the parameters of the small CLIP model, the specific loss formula is shown as follows:
$\mathcal{L}_{\text{final}}=\mathcal{L}_{\text{stu}}+\mathcal{L}_{\text{align}}.$ (9)
@@
Following (Zhou et al. 2022a) (Khattak et al. 2023a) (Khattak et al. 2023b) (Li et al. 2024b), we divide the training and testing datasets into base and novel classes. Our teacher model, pre-trained via the PromptSRC approach (Khattak et al. 2023b), guides the training of student model using an unlabeled set. Post-distillation, we assess the students’ performance on both class types against the test set, serving as a measure of methodological generalization within the dataset. As shown in Table 1, we compare the performance of our proposed ComKD-CLIP with recent state-of-the-art methods including CLIP (Radford et al. 2021), CoOp (Zhou et al. 2022b), CoCoOp (Zhou et al. 2022a), MaPLe (Khattak et al. 2023a), PromptSRC (Khattak et al. 2023b), PromptKD (Li et al. 2024b) on 11 recognition datasets. In comparison with these state-of-the-art works, ComKD-CLIP shows highly competitive results in all 11 datasets.
@@
Similar to PromptKD (Li et al. 2024b), our teacher model undergoes pre-training on ImageNet (Deng et al. 2009). Subsequently, we utilize the training set of unlabeled target datasets to train the student model. Their performance is then evaluated on the test set post-training, with no data-specific fine-tuning applied. We compare our cross-dataset performance with previous methods in Table 2, our proposed ComKD-CLIP outperforms some state-of-the-art methods on 8 out of 10 datasets, achieving an average improvement of 0.74% over previous methods.
@@
Table 2: The performance of ComKD-CLIP with some state-of-the-art mothods in a cross-dataset benchmark. Utilizing our pipeline, we conduct unsupervised Aligned distillation with unlabeled domain data in a transductive setting. The source model is pretrained on ImageNet. ”ZSL” indicates that the evaluation is conducted in a Zero-Shot Learning setting. ComKD-CLIP outperforms on 8 out of 10 datasets.
@@
Table 3: The performance of ComKD-CLIP within domain generalization contexts. The method is trained on ImageNet and subsequently evaluated on datasets with domain shifts.
@@
To begin with, we try to explore the effectiveness of IFAlign module and EduAttention module. Specifically, we remove the IFAlign module and EduAttention module from ComKD-CLIP respectively and test the corresponding experimental results. As shown in Table 4, removing IFAlign module will greatly reduce the performance of ComKD-CLIP. Similarly, removing the EduAttention module also decrease the model’s performance. It is worth noting that when both the IFAlign module and the EduAttention module are removed, ComKD-CLIP has the worst performance. These experimental results fully prove that IFAlign module and EduAttention module can effectively prompt the student model to absorb the teacher model’s knowledge on how to extract image features and integrate text-image feature.
@@
Subsequently, we try to use different alignment methods in IFAlign to find the best alignment strategy. We respectively align image features by mean ($\mathcal{L}_{\text{align\_mean}}$), variance ($\mathcal{L}_{\text{align\_var}}$), and mean add variance ($\mathcal{L}_{\text{align}}$). The experimental results are shown in Table 5. It can be clearly seen that ComKD-CLIP has the best performance when we use the mean add variance ($\mathcal{L}_{\text{align}}$) to align image features in IFAlign. This mainly because that mean add variance can jointly promote the alignment from the perspectives of center position and discreteness, thereby helping the student model to absorb the teacher model’s knowledge on how to extract image features.
@@
Finally, we try to explore the best strategy to refine the knowledge distilled by IFAlign module and EduAttention module. Specifically, we use KL divergence, L1 and MSE methods to refine the distilled knowledge. The corresponding performance of ComKD-CLIP is shown in Table 6. It can be seen that KL divergence can better refine the distilled knowledge and make ComKD-CLIP perform best. This may be because KL divergence can make the logits of the student model better approximate the logits of the teacher model, thus more effectively refining the knowledge distilled by the student model.